4 squirrels were found to have an average weight of 9.3 ounces with a sample standard deviation is 1.1. Find the 95% confidence interval of the true mean weight.




Answer by Boreal(15235)

The half interval is t(0.975, df=3)*s/sqrt(n)
=3.182*1.1/2=1.75
The interval is mean +/- half interval, or (7.55, 11.05) units ounces